# Improved Belief Propagation Decoding on Surface Codes with High Accuracy and Low Latency

### v1

@@ 3.1.1 Update direction smoothing
Momentum-BP can be implemented in both parallel and serial scheduling. Here, we provide the latter in algorithm 3.1.1, where the initialization and check node update remain the same as in traditional BP. Our approach computes the a posteriori probability before performs the vertical update. The posterior update rule of Momentum-BP is shown as Eq. (16) and Eq. (17) where $g_{v_{j},e}^{(t)}$ are the accumulated momentum term and $Q_{v_{j},e}^{(t)}$ are the a posteriori probability for the $j$-th variable node regarding the Pauli error $e$ in the $t$-th iteration. Notice that we transformed Eq. (8), so that the accumulated term is the posterior value from the previous iteration. In the variable node update step, we subtract the original message sent from the check node rather than the momentum-smoothed message to ensure stability and prevent overcompensation.
{algorithm}
[!ht] Momentum-BP under Serial Scheduling for Quaternary Log Domain
Input: Quaternary parity-check matrix $\boldsymbol{H}$ of $M\times N$, syndrome vector $\boldsymbol{s}=(1,-1)^{M}$, a priori probability vectors $\{p_{v_{j}}^{X},p_{v_{j}}^{Y},p_{v_{j}}^{Z}\}_{j=1}^{N}$, max iterations $iter_{max}$, learning rate $\alpha=1$, and momentum rate $\gamma\geq 0$.
@@
if $\langle H,\hat{E}\rangle=s$, return "converge" and halt.
else if $t=iter_{max}$, return "fail" and halt.
@@
The properties of Momentum-BP can be explained from two perspectives. (1) Since it takes into account message updates from previous iterations, the method smooths the update direction in each iteration, preventing the corresponding energy function from oscillating between two points and gradually converging to the global minimum between them. (2) Node-wise momentum updates gradually break the symmetry of messages in the trapping sets shown in Fig. 2, allowing BP to escape from these trapping sets. The trapping set simulation in Section 4.1 will support those explanations.
3.1.2 Adaptive step size
@@
3.2 Dynamic Initial probability
BP decoding and gradient descent still have a fundamental difference, which limits the performance of algorithms in Section 3.1. Each update in gradient descent is based on the previous step’s results, whereas in BP decoding, each calculation of a posteriori probabilities accumulates the same a priori probabilities. Given the objective function as shown in Eq. (2), for simplicity using the binary domain for instance, BP calculates the LLR values as
@@ 3.2 Dynamic Initial probability
Existing methods for altering initial probabilities include re-initialization using information obtained from previous iterations [20, 21, 30], as well as detecting oscillations during the decoding process and modifying initial values accordingly [31]. However, they either require multiple outer loops or manually set algorithm parameters. This subsection introduces two methods for dynamically updating initial values, utilizing a posteriori probabilities from previous iterations to accelerate the convergence of BP decoding with almost no additional overhead.
3.2.1 EWA Initialization
The first method employs a transformed exponential weighted averaging, incorporating components of all previous a posteriori probabilities into the a priori probabilities with a decaying factor. EWAInit-BP under parallel scheduling is shown in algorithm 3.2.1.
{algorithm}
[!ht] EWAInit-BP under Parallel Scheduling for Quaternary Log Domain
@@ 3.2.1 EWA Initialization
if $\langle H,\hat{E}\rangle=s$, return "converge" and halt.
else if $t=iter_{max}$, return "fail" and halt.
@@
In fact, combining Eq. (27) and Eq. (8), EWAInit-BP exhibits similar characteristics to Momentum-BP, namely the exponentially decaying components of previous iterations. The difference is that the former aggressively changes the "prior" probability of each variable node in each iteration. Eq. (27) still retains a portion of the input prior probabilities. In our experiments, we tried replacing this term with the prior probabilities from the previous iteration, but the algorithm’s performance did not significantly change.
3.2.2 Adaptive Initialization
The second method is similar to the AdaGrad approach used in message updates, except that it updates the prior probabilities, shown as
$g_{v_{j},e}^{(t)}$ $=Q_{v_{j},e}^{(t-1)}-Q_{v_{j},e}^{(0)}$ (28)
$G_{v_{j},e}^{(t)}$ $=G_{v_{t},e}^{(t-1)}+(g_{v_{j},e}^{(t)})^{2}$ (29)
$\Pi_{v_{j},e}^{(t)}$ $=Q_{v_{j},e}^{(0)}+\frac{\alpha g_{v_{j},e}^{(t)}}{\sqrt{G_{v_{j},e}^{(t)}+\epsilon}}$ (30)
The core concept involves calculating the difference between the a posteriori probabilities from the previous iteration and the initial a priori probability, and updating this difference with a decayed learning rate.
3.3 Complexity
Our proposed algorithms involve the same message passing between variable nodes and check nodes as traditional BP. The main difference lies in the additional update steps, which in Momentum-BP and AdaGrad-BP consist of recalculating the a posteriori probabilities, and in EWAInit-BP and AdaInit-BP involve recalculating the initial values. These additional steps only involve simple arithmetic operations and can be executed locally at each node, without requiring extra iterations, loops, or global communication, thus not introducing any significant overhead. Therefore, the overall time complexity of the proposed algorithms remains $O(j\tau N)$, where $j$ is the average degree of the Tanner graph, and $\tau$ is the number of iterations.
@@ 3.3 Complexity
4 Simulation Results
In all simulations, the criterion for successful decoding is that the error estimate satisfies $\hat{E}\in E\mathcal{S}$, rather than $\hat{E}=E$ as in classical information. All cases not meeting the above condition are referred to as logical errors in this paper. Herein, a detected error refers to $\hat{E}E$ being anti-commutative with any one stabilizer $S\in\mathcal{S}$; an undetected error implies $\hat{E}E$ commutes with all stabilizers, i.e., $\hat{E}E\in N(\mathcal{S})\backslash\mathcal{S}$, where the result of error correction is a logical operator. Each data point on the experimental curves represents the mean of at least $10^{4}$ simulations.
Our simulations employ a depolarizing channel, where each qubit has an equal probability $p/3$ of experiencing X, Y, or Z Pauli errors, and a probability of $1-p$ of no error occurring. Additionally, for ease of comparison, we adopt the simple noise model which does not consider measurement errors or error propagation in quantum circuits.
4.1 Breaking the Trapping Sets
We first compare the performance of improved BP on the trapping set shown in Fig. 2. In this example, qubit 1 and qubit 2 experience X errors, resulting in an error syndrome of [1,0,1,0]. Fig. 3 shows the comparison between the traditional LLR-BP4, Momentum-BP and AdaGrad-BP in Section 3.1 under parallel scheduling, with the maximum iteration set to 50. The hard decision result of the traditional BP oscillates continuously between all X and all I, while the improved BP shows significantly reduced oscillation amplitude and converges within 12 iterations.
@@ 4.1 Breaking the Trapping Sets
Algorithm Parallel Scheduling Serial Scheduling
@@
MBP [25] ($\alpha=0.6$) 2 iterations
BP-OTS [31] (Different T) T+2 iterations T+2 iterations
Momentum-BP($\alpha=0.6$) 12 iterations 6 iterations
AdaGrad-BP($\alpha=5$) 12 iterations 2 iterations
EWAInit-BP($\alpha=0.6$) 2 iterations
A more detailed comparison of the different algorithms proposed in this paper is presented in Table. 1. The parameters used for the algorithms are defaults and were not specially optimized for this example. MBP [25] converges at the same fastest speed under serial scheduling but oscillates under parallel scheduling. BP-OTS [31] shows little dependency on scheduling method but its convergence depends on the manually set oscillation period $T$, which may vary in different error correction scenarios. The algorithms proposed in Section 3.1, Momentum-BP and AdaGrad-BP, both converge under parallel scheduling. Besides, AdaGrad-BP and EWAInit-BP with adaptive a priori probabilities converge at the fastest speed under serial scheduling.
4.2 BP Optimization on Topological codes
In this subsection, we compare the performance of the various BP optimization proposed in this paper. Subsequently, we present comparisons of our algorithm with other BP improvements without post-processing on Toric codes, Surface codes, and XZZX Surface codes. For all decoding algorithms, the maximum number of iterations $iter_{max}$ across all code lengths is set to 150 to ensure linear asymptotic complexity.
4.2.1 Comparison of our BP optimization methods
Figure 4: Performance of various BP optimization methods presented in this paper, demonstrated using the $L=8$ toric code as an example. Each method uses default parameters and has not been optimized for this specific case. Among these, the method for optimizing initial values, EWAInit-BP, exhibits the best performance and completely eliminates the error floor.
A comparison of all methods proposed in this paper is illustrated in Fig. 4. The two methods for optimizing message update rules, Momentum-BP and AdaGrad-BP, exhibit some improvements over traditional BP at low physical error rates, while their performance gradually converging to that of traditional BP as the physical error rate increases. The two methods for optimizing initial values, AdaInit-BP and EWAInit-BP, demonstrate improvements ranging from 1 to 4 orders of magnitude over traditional BP, and excellently eliminate the error floor. However, the performance of algorithms that combine these two ideas is not satisfactory, because both approaches inherently consider the historical a posteriori probabilities but from different positions within the BP algorithm. This two methods are essentially addressing the same problem—incorporating past information to refine current updates. As a result, simply merging them can lead to redundancy and overcompensation, which negatively impacts the overall performance.
4.2.2 Toric code
We construct the $[[2n^{2},1,n]]$ toric code based on its properties as a hypergraph product of an $[n,1]$ repetition code. The toric code features periodic boundaries, hence all its stabilizers have a weight of four, and any error syndromes on it exhibit translational symmetry [50], which is particularly conducive to BP and neural network decoding. We compare the performance of our EWAInit-BP and other existing methods under both parallel and serial scheduling as shown in Fig. 5(a) and Fig. 6(a), respectively.
@@ 4.2.2 Toric code
Figure 5: Performance of LLRBP4, MBP, and EWAInit-BP under parallel scheduling on toric codes with lattice sizes of 4, 6, 8. Only EWAInit-BP demonstrates a reduction in the logical error rate as the lattice size increases, with the intersection of error rates occurring at 13.6%.
@@
Figure 6: Performance of MBP, BP-OTS, and EWAInit-BP under serial scheduling on toric codes with L = 4, 6, 8. We choose $T=9$ and $C=20$ for BP-OTS. EWAInit-BP outperforms other BP improvements, with the intersection of logical error rates at 15%.
For the parallel schedule, compared to the classic logarithmic domain quaternary BP and its enhancement, MBP, our EWAInit-BP is the only method that achieves a decrease in logical error rates as the lattice size increases, with significant improvements in logical error rates at the same lattice size. [^1] Our implementation of BP-OTS in the quaternary domain with parallel scheduling did not achieve the performance reported in the original paper, necessitating further investigation into this discrepancy. The intersection of the error rate curves for these lattice sizes is approximately 13.6%. Additionally, the average number of iterations required for EWAInit-BP is significantly lower than the other two methods, converging in fewer than 10 iterations mostly. Parallel scheduling of BP algorithms holds the potential to achieve constant complexity through hardware-based parallel computation, while EWAInit-BP ensures both the error correction capabilities and speed in decoding.
@@
4.2.3 Planar surface code
@@
Figure 7: Performance of LLRBP4, MBP, and EWAInit-BP (EWAInit-OTS) on planar surface codes with L = 5, 7, 9. The intersection of logical error rates for EWAInit-BP under serial scheduling is approximately 10%.
@@
Under serial scheduling, EWAInit-BP performs better than MBP at $L=5$ and $L=7$; however, its performance at $L=9$ is not satisfying. In fact, the decoding failure rate of EWAInit-BP increases with higher code lengths, thus not fully demonstrating the error-correcting capabilities of the code. This is because, despite the dynamic updating of a priori probabilities, the method still partly relies on the initial input priors. We incorporate the OTS method as a "post-processing" step for EWAInit, resetting the a priori probabilities of some nodes every few iterations. We choose parameter T=5 and C=20. This EWAInit-OTS method, while maintaining a time complexity of $O(n)$, shows improvements in convergence rate and logical error rates on the surface code compared to EWAInit-BP.
4.2.4 XZZX Surface code
@@
Figure 8: Performance of LLRBP4, MBP, and EWAInit-BP (EWAInit-OTS) on XZZX surface codes with L = 5, 7, 9 under serial scheduling. The intersection of logical error rates for EWAInit-BP is approximately 12.4%.
@@
Figure 9: Performance of LLRBP4 and EWAInit-BP on the XZZX surface code with L = 5, 7, 9 under parallel scheduling, evaluated under conditions where Pauli Z errors are more prevalent in the channel. The intersection of LERs for EWAInit-BP is approximately 10%.
@@
5 Summary
We analyzed the reasons for BP’s poor performance on surface codes, namely trapping sets formed by 8-cycles and symmetric degeneracy. In these trapping sets, the probability vectors in each iteration exhibit periodic oscillations. We proposed several methods to mitigate this issue.
Since the a posteriori probability update of BP can be interpreted as the gradient optimization of an energy function, we borrowed optimization strategies from gradient descent in machine learning, proposing Momentum-BP and AdaGrad-BP. These mechanisms aim to smooth the update direction and reduce the amplitude of oscillations, respectively. Experiments on trapping sets intuitively demonstrated the effectiveness of these algorithms.
From the theoretical derivation of the BP algorithm, we identified the importance of a priori probabilities and transformed the message update smoothing into initial values adapting, proposing EWAInit-BP. This algorithm exhibits aggressive exploration and outperforms state-of-the-art algorithms without OSD post-processing on the Toric code, Surface code, and XZZX Surface code. Specifically, EWAInit-BP can demonstrate error correction capabilities even under parallel scheduling, with logical error rates decreasing as code length increases. Notably, all our proposed methods do not introduce any additional factors to the time complexity of BP. Thus, EWAInit-BP can be implemented with $O(1)$ complexity in fully parallel while maintaining decoding accuracy.
Parameterized BP algorithms involve a trade-off between convergence and the occurrence of logical operators. Our methods primarily focus on ensuring convergence, as the main issue with BP on surface codes lies in achieving this, which subsequently leads to an increased rate of logical operators. Therefore, it is challenging to apply our improvements directly to general quantum error-correcting codes. A potential research direction is to use AI methods to explore this trade-off and find the optimal balance for the lowest error rate.
Additionally, our simulations only considered Pauli errors on qubits and did not account for measurement errors, error propagation in circuits, etc. Future work needs to adapt BP to more challenging noise models, potentially requiring specialized optimizations.

### v2

@@ III.1.1 Update smoothing
Momentum-BP can be implemented under both parallel and serial scheduling. Here, we provide the latter in algorithm 1, where the initialization and check node update remain the same as in traditional BP. Our approach computes the a posteriori probability before performs the vertical update. The posterior update rule of Momentum-BP is shown as Eq. (16) and Eq. (17) where $g_{v_{j},e}^{(t)}$ are the accumulated momentum term and $Q_{v_{j},e}^{(t)}$ are the a posteriori probability for the $j$-th variable node regarding the Pauli error $e$ in the $t$-th iteration. Notice that we transformed Eq. (8), so that the accumulated term is the posterior value from the previous iteration. In the variable node update step, we subtract the original message sent from the check node rather than the momentum-smoothed message to ensure stability and prevent overcompensation.
Algorithm 1 Momentum-BP under Serial Scheduling over GF(4)
Input: Parity-check matrix $\boldsymbol{H}$ of $M\times N$, syndrome vector $\boldsymbol{s}=(1,-1)^{M}$, a priori probability vectors $\{p_{v_{j}}^{X},p_{v_{j}}^{Y},p_{v_{j}}^{Z}\}_{j=1}^{N}$, max iterations $iter_{max}$, learning rate $\alpha=1$, and momentum rate $\gamma\geq 0$.
@@
if $\langle H,\hat{E}\rangle=s$, return ”converge” and halt.
else if $t=iter_{max}$, return ”fail” and halt.
@@
The properties of Momentum-BP can be explained from two perspectives. (1) Since it takes into account message updates from previous iterations, the method smooths the update direction in each iteration, preventing the corresponding energy function from oscillating between two points and gradually converging to the global minimum between them. (2) Node-wise momentum updates gradually break the symmetry of messages in the trapping sets shown in Fig. 2, allowing BP to escape from these trapping sets. The trapping set simulation in Section IV.1 will support those explanations.
III.1.2 Adaptive step size
@@
III.2 EWA Initialization of a priori probability
BP decoding and gradient descent still have a fundamental difference, which limits the performance of algorithms in Section III.1. Each update in gradient descent is based on the previous step’s results, whereas in BP decoding, each calculation of a posteriori probabilities accumulates the same a priori probabilities. Given the objective function as shown in Eq. (2), for simplicity using the binary domain for instance, BP calculates the LLR values as
@@ III.2 EWA Initialization of a priori probability
Existing methods for altering initial probabilities include re-initialization using information obtained from previous iterations [22, 23, 31], as well as detecting oscillations during the decoding process and modifying initial values accordingly [32]. However, they either require multiple outer loops or manually set algorithm parameters. This subsection introduces a method for dynamically updating a priori probabilities, employing a transformed exponential weighted averaging, incorporating components of all previous a posteriori probabilities into the a priori probabilities with a decaying factor. Hence, we call this method the EWAInit-BP. EWAInit-BP under parallel scheduling is shown in algorithm 2.
Algorithm 2 EWAInit-BP under Parallel Scheduling over GF(4)
@@
if $\langle H,\hat{E}\rangle=s$, return ”converge” and halt.
else if $t=iter_{max}$, return ”fail” and halt.
@@
In fact, combining Eq. (27) and Eq. (8), EWAInit-BP exhibits similar characteristics to Momentum-BP, namely the exponentially decaying components of previous iterations. The difference is that the former aggressively changes the ”prior” probability of each variable node in each iteration. Eq. (27) still retains a portion of the input prior probabilities. In our experiments, we tried replacing this term with the prior probabilities from the previous iteration, but the algorithm’s performance did not significantly change.
III.3 Complexity
Our proposed algorithms involve the same message passing between variable nodes and check nodes as traditional BP. The main difference lies in the additional update steps, which in Momentum-BP and AdaGrad-BP consist of recalculating the a posteriori probabilities, and in EWAInit-BP involves recalculating the a priori probabilities. These additional steps only involve simple arithmetic operations and can be executed locally at each node, without requiring extra iterations, loops, or global communication, thus not introducing any significant overhead. Therefore, the overall time complexity of the proposed algorithms remains $O(j\tau N)$, where $j$ is the average degree of the Tanner graph, and $\tau$ is the number of iterations.
@@ III.3 Complexity
IV Simulation Results
In all simulations, the criterion for successful decoding is that the error estimate satisfies $\hat{E}\in E\mathcal{S}$, rather than $\hat{E}=E$ as in classical information. All cases not meeting the above condition are referred to as logical errors in this article. Herein, a detected error refers to $\hat{E}E$ being anti-commutative with any one stabilizer $S\in\mathcal{S}$; an undetected error implies $\hat{E}E$ commutes with all stabilizers, i.e., $\hat{E}E\in N(\mathcal{S})\backslash\mathcal{S}$, where the result of error correction is a logical operator. Each data point on the experimental curves represents the mean of at least $10^{4}$ simulations.
Our simulations employ a depolarizing channel, where each qubit has an equal probability $p/3$ of experiencing Pauli $X$, $Y$, or $Z$ errors, and a probability of $1-p$ of no error occurring. Additionally, for ease of comparison, we adopt the simple noise model which does not consider measurement errors or error propagation in quantum circuits.
IV.1 Breaking the Trapping Sets
@@
We first compare the performance of improved BP on the trapping set shown in Fig. 2. In this example, qubit 1 and qubit 2 experience X errors, resulting in an error syndrome of $[1,0,1,0]$. Fig. 3 shows the comparison between the traditional LLR-BP4, Momentum-BP and AdaGrad-BP in Section III.1 under parallel scheduling, with the maximum iteration set to 50. The hard decision result of the traditional BP oscillates continuously between all $X$ and all $I$, while improved BP algorithms show significantly reduced oscillation amplitude and converge within 12 iterations.
A more detailed comparison of the different algorithms proposed in this article is presented in Table. 1. The parameters used for these algorithms are defaults and were not specially optimized for this example. MBP [20] converges at the same fastest speed under serial scheduling but oscillates under parallel scheduling. BP-OTS [32] shows little dependency on scheduling method but its convergence depends on the manually set oscillation period $T$, which may vary in different error correction scenarios. Algorithms proposed in Section III.1, Momentum-BP and AdaGrad-BP, both converge under parallel scheduling. Besides, AdaGrad-BP and EWAInit-BP with adaptive a priori probabilities converge at the fastest speed under serial scheduling.
@@
Algorithm  [^1] The parameters used for each algorithm are: MBP ($\alpha=0.6$), Momentum-BP ($\alpha=0.6$), AdaGrad-BP ($\alpha=5$), EWAInit-BP ($\alpha=0.6$) Parallel Scheduling Serial Scheduling
@@
MBP [20] 2
BP-OTS [32] T+2 T+2
EWAInit-BP 2
Momentum-BP 12 6
AdaGrad-BP 12 2
IV.2 Improved BP on Topological codes
In this subsection, we compare the performance of various BP improvements proposed in this article. Subsequently, we present comparisons of our algorithm with other BP improvements without post-processing for toric codes, planar surface codes, and XZZX surface codes. For all decoding algorithms, the maximum number of iterations $iter_{max}$ across all code lengths is set to 150 to ensure linear asymptotic complexity.
IV.2.1 Comparison of our BP optimization methods
Figure 4: Performance of various BP optimization methods presented in this article under serial scheduling, demonstrated using the $L=8$ toric code as an example. Each method uses default parameters. Among these, the method for optimizing initial values, EWAInit-BP, exhibits the best performance and completely eliminates the error floor.
A comparison of all methods proposed in this article is illustrated in Fig. 4. The two methods for optimizing message update rules, Momentum-BP and AdaGrad-BP, exhibit some improvements over traditional BP at low physical error rates, while their performance gradually converging to that of traditional BP as the physical error rate increases. The method for optimizing initial values, EWAInit-BP, demonstrate improvements ranging from 1 to 3 orders of magnitude over traditional BP, and excellently eliminate the error floor. However, the performance of algorithms that combine these two ideas is not satisfactory, because both approaches inherently consider the historical a posteriori probabilities but from different positions within the BP algorithm. This two methods are essentially addressing the same problem—incorporating past information to refine current updates. As a result, simply merging them can lead to redundancy and overcompensation, which negatively impacts the overall performance.
IV.2.2 Toric code
We construct the $[[2n^{2},1,n]]$ toric code based on its properties as a hypergraph product of an $[n,1]$ repetition code. The toric code features periodic boundaries, hence all its stabilizers have a weight of four, and any error syndromes on it exhibit translational symmetry [16], which is particularly conducive to BP and neural network decoding. We compare the performance of our EWAInit-BP and other existing methods under both parallel and serial scheduling as shown in Fig. 5 and Fig. 6, respectively.
@@ IV.2.2 Toric code
Figure 5: Performance of LLRBP4, MBP, and EWAInit-BP under parallel scheduling for toric codes with L = {4, 6, 8}. Only EWAInit-BP find a reduction in the logical error rate as the lattice size increases, with the intersection of error rates occurring at 13.6%.
@@
Figure 6: Performance of MBP, BP-OTS, and EWAInit-BP under serial scheduling for toric codes with L = {4, 6, 8}. We choose $T=9$ and $C=20$ for BP-OTS. EWAInit-BP outperforms other BP improvements, with the intersection of logical error rates at 15%.
For the parallel schedule  [^2] Our implementation of BP-OTS over GF(4) under parallel scheduling did not achieve the performance reported in the original paper, necessitating further investigation into this discrepancy., compared to the classic BP over GF(4) with LLR message and its enhancement, MBP, our EWAInit-BP is the only method that achieves a decrease in logical error rates as the lattice size increases, with significant improvements in logical error rates at the same lattice size. The intersection of the error rate curves for these lattice sizes is approximately 13.6%. Additionally, the average number of iterations required for EWAInit-BP is significantly lower than the other two methods, converging in fewer than 10 iterations mostly. Parallel scheduling of BP algorithms holds the potential to achieve constant complexity through hardware-based parallel computation, while EWAInit-BP ensures both the error correction capabilities and speed in decoding.
@@
IV.2.3 Planar surface code
@@
Figure 7: Performance of LLRBP4, MBP, and EWAInit-BP (EWAInit-OTS) for planar surface codes with L = 5, 7, 9. The intersection of logical error rates for EWAInit-BP under serial scheduling is approximately 10%.
@@
Under serial scheduling, EWAInit-BP performs better than MBP at $L=5$ and $L=7$; however, its performance at $L=9$ is not satisfying. In fact, the decoding failure rate of EWAInit-BP increases with higher code lengths, thus not fully demonstrating the error-correcting capabilities of the code. This is because, despite the dynamic updating of a priori probabilities, the method still partly relies on the initial input priors. We incorporate the OTS method as a ”post-processing” step for EWAInit, resetting the a priori probabilities of some nodes every few iterations, with the hyperparameters T=5 and C=20. This EWAInit-OTS method, while maintaining a time complexity of $O(n)$, shows improvements in convergence rate and logical error rates on the surface code compared to EWAInit-BP and other methods.
The patterns of average iterations of different algorithms for the planar surface code and the XZZX surface code are similar to those observed in the toric code and exhibit the same trends as the logical error rates. Therefore, we have omitted this data.
IV.2.4 XZZX Surface code
@@
Figure 8: Performance of LLRBP4, MBP, and EWAInit-BP (EWAInit-OTS) for XZZX surface codes with L = 5, 7, 9 under serial scheduling. The intersection of logical error rates for EWAInit-BP is approximately 12.4%.
@@
Figure 9: Performance of LLRBP4 and EWAInit-BP for the XZZX surface code with L = 5, 7, 9 under parallel scheduling, evaluated under conditions where Pauli Z errors are more prevalent in the channel. The intersection of LERs for EWAInit-BP is approximately 10%.
@@
V Conclusion
In this article, we proposed three improved BP algorithms which are Momentum-BP, AdaGrad-BP and EWAInit-BP.
Since the a posteriori probability update of BP can be interpreted as the gradient optimization of an energy function, we borrowed optimization strategies from gradient descent in machine learning, proposing Momentum-BP and AdaGrad-BP. These mechanisms aim to smooth the update process and reduce the amplitude of oscillations, respectively. Experiments on trapping sets intuitively demonstrated the effectiveness of these algorithms.
From the theoretical derivation of the message update rules of BP algorithm, we realized the importance of a priori probabilities and transformed the message update smoothing into initial values adapting, proposing EWAInit-BP. This algorithm exhibits aggressive exploration and outperforms other algorithms without OSD post-processing on the toric code, planar surface code, and XZZX surface code. Specifically, EWAInit-BP can demonstrate error correction capabilities even under parallel scheduling, with logical error rates decreasing as code length increases. Notably, all our proposed methods do not introduce any additional factors to the time complexity of BP. Thus, EWAInit-BP can be implemented with $O(1)$ complexity in fully parallel while maintaining decoding accuracy.
However, the aggressive exploration and convergence capabilities of EWAInit increase the probability of generating logical operators after error recovery. Therefore, directly applying this method to other error-correcting codes presents certain challenges. A potential research direction is to identify situations where logical operators may occur. Besides, our simulations only considered Pauli errors on qubits and did not account for measurement errors, error propagation in circuits, etc. Future work needs to adapt BP to more challenging noise models, potentially requiring specialized optimizations.
Acknowledgements.
We would like to express our sincere gratitude to Zhengzhong Yi for his in-depth discussions on the principles of BP algorithms and meticulous feedback on the manuscript. We are also thankful to Zhipeng Liang for his assistance with the experimental code framework. Finally, we extend our appreciation to Prof. Xuan Wang for his insightful discussions and guidance on the gradient approach.